# Anti-bosonic current in thermal spin pumping

### v1

@@ Abstract
We investigate thermal spin pumping in gadolinium iron garnet, focusing on dynamics of antiferromagnetic magnon and its impact on spin and heat transport. Antiferromagnet has right-handed and left-handed magnon modes. Using a two-magnetic-lattice model, we derive temperature-dependent spin-wave spectra and thermal spin pumping constants, revealing distinct transport regimes governed by competing contribution of magnon modes. We show that competition of both modes induces a compensation temperature in thermally generated spin current. This theoretical framework deepens the understanding of spin-caloritronic effects in ferrimagnets, offering new perspectives for spintronic device optimization.
Introduction. The exploration of fundamental particles has driven significant advancements in our understanding of quantum mechanics and condensed matter physics. Among these particles, the study of antiferromagnetic magnons, i.e., the quanta of spin waves in antiferromagnetic systems, plays a pivotal role in various theoretical frameworks, including quantum field theory and spintronics. Magnons are essential for understanding phenomena such as spin transport and the manipulation of spin states, which are crucial in quantum information science and spintronics Lebrun et al. (2020); Cornelissen et al. (2015). Recent advancements have demonstrated that antiferromagnets can be explored as promising systems for spintronic applications, including quantum computing and communication technologies Zhou et al. (2022); Poelchen et al. (2023).
Building on the concept of magnons, researchers have introduced the notion of magnons with different handedness, which are spin-wave excitations that can precess in opposite directions, depending on the underlying magnetic structure. In particular, the antiferromagnetic magnon in systems with alternating spin orientations exhibits different modes of propagation, including left-handed and right-handed modes Kanj et al. (2023), as illustrated in Fig. 1. Detecting antiferromagnetic magnons with distinct handedness represents an emerging frontier in condensed matter physics, with techniques such as thermal spin pumping offering promising potential Siddiqui et al. (2021). Thermal spin pumping utilizes thermally excited spin waves to generate electrical signals, providing an indirect method to probe magnonic properties. By integrating thermal spin pumping with advanced detection techniques, researchers could pave the way for groundbreaking discoveries in quantum magnonics and spin-based computing Ortiz et al. (2021). The coupling of thermal spin pumping with magnons could facilitate the manipulation and detection of spin waves, particularly those with unique handedness, thereby enhancing their applicability for future technological developments Holzmann et al. (2022).
Here, we proposed that the distinct chirality of the antiferromagnetic magnon modes can be conceptualized as bosonic quasiparticles analogous to how holes in semiconductor systems are considered antifermionic quasiparticles. This novel perspective allows us to view the handedness of magnons as a form of antiboson, thereby enriching the theoretical framework of magnonics Lachance-Quirion et al. (2019). The potential applications of these distinct modes also hold promise for innovative approaches in quantum magnonics and the manipulation of information via spintronic devices Weiss et al. (2023); Bai et al. (2015).
@@
Rare earth iron garnets (RIGs) are particularly well-suited for studying antiferromagnetic magnons due to their unique magnetic properties and exceptionally low magnetic damping A. Seo et al. (2024). The intricate magnetic structure of RIGs—including three magnetic sublattices (tetrahedral Fe3+ sites, octahedral Fe3+ sites, and dodecahedral rare earth R3+ sites)—generates a rich spin wave spectrum with multiple branches Barker and Bauer (2016); Huang et al. (2024). Tuning the rare earth ion composition allows for precise engineering of the spin wave spectrum, potentially facilitating the detection of magnons with distinct handedness Zhang et al. (2023). Furthermore, RIGs, particularly yttrium iron garnet (YIG), exhibit exceptionally low magnetic damping, rendering them ideal for studying spin wave dynamics and the integration of spintronic devices Cornelissen et al. (2015); Li et al. (2020). The presence of rare earth ions in RIGs introduces strong spin-orbit coupling effects, which may induce phenomena such as the inverse Faraday effect and the Dzyaloshinskii-Moriya interaction, critical for the manipulation and detection of spin waves with distinct handedness Cheng et al. (2016).
YIG thermal spin pumping. In thermal spin pumping, spin $J_{s}$ and heat $J_{Q}$ currents are transferred from a magnetic material, such as rare-earth iron garnet, to a non-magnetic metal, such as Pt. This spin current is then converted into an electrical signal via the inverse spin Hall effect. At the interface, $J_{s}$ and $J_{Q}$ coupled and can be expressed through a linear matrix relation involving the driving forces: spin accumulation $\bm{\mu}_{s}$ and temperature difference $\Delta T$, analogous to the thermoelectric effect:
$\left[\begin{array}[]{c}J_{s}\\ J_{Q}\end{array}\right]=\left[\begin{array}[]{cc}\mathcal{L}_{0}&\mathcal{L}_{1}\\ \mathcal{L}_{1}&\mathcal{L}_{2}\end{array}\right]\left[\begin{array}[]{c}-\hat{\textbf{m}}\cdot\bm{\mu}_{s}\\ -\Delta T/T\end{array}\right].$ (1)
Here, $\hat{\textbf{m}}=z\hat{z}S_{z}/|S_{z}|$ is the average direction of the magnetization. $\mathcal{L}_{i}$ is analogous to thermoelectric transport coefficients. $\mathcal{L}_{0}$ represents the interface spin injection conductance, while $\mathcal{L}_{2}$ denotes the heat conductance. The ratios $\mathcal{L}_{1}/\mathcal{L}_{0}$ and $\mathcal{L}_{1}/(\mathcal{L}_{0}T)$ correspond to the spin Peltier and spin Seebeck coefficients, respectively Cahaya et al. (2014).
@@
$J_{\mathrm{fl}}=-\gamma\left<\textbf{S}\times\textbf{h}^{\prime}\right>\cdot\hat{z}$ is spin backflow due to thermal noise. $J_{\mathrm{fl}}=-J_{\mathrm{sp}}(T-\Delta T)/T$ has been shown to balances $J_{\mathrm{sp}}$ when there is no temperature difference across interface Xiao et al. (2010).
@@
where $G$ is the spin-mixing conductanceJiao and Bauer (2013).
@@
[…]
@@
$\mathcal{L}_{0}=$ $2GS_{z}\int_{\omega_{0}}^{\infty}\frac{\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}=\frac{GS_{z}}{2}\left(\frac{k_{B}T}{\pi D}\right)^{\frac{3}{2}}\mathrm{Li}_{\frac{3}{2}}\left(e^{-\beta\omega_{0}}\right),$ (7)
@@
$=2GS_{z}\int_{0}^{\infty}\frac{\omega\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}=\frac{3GS_{z}\left(k_{B}T\right)^{\frac{5}{2}}}{4\left(\pi D\right)^{\frac{3}{2}}}\mathrm{Li}_{\frac{5}{2}}\left(e^{-\beta\omega_{0}}\right).$ (8)
In YIG, the magnon gap $\omega_{0}=\gamma H$ can be non zero due to Zeeman effect Rodrigue et al. (1960). Since $\lim_{x\to 0}\mathrm{Li}_{n}(x)\simeq x$, the magnetic field $H$ exponentially suppress thermal spin pumping Kikkawa et al. (2015) $\mathcal{L}_{1}\propto e^{-\gamma H/k_{B}T}.$ The magnetization of YIG originates from five Fe3+ ions distributed across two magnetic lattices. Two of these ions occupy tetrahedral Fe3+ sites, while three are in octahedral sites. Due to their antiferromagnetic coupling, only one Fe3+ spin contributes Néel et al. (1964), as illustrated in Fig. 2a. When yttrium in YIG is substituted by lanthanides, R3+ occupies the dodecahedral site, forming an additional magnetic lattice. This substitution leads to a more complex spin-wave spectrum compared to that of YIG. While significant computational efforts are underway to unravel these spectra Wang et al. (2020); Li et al. (2022), determining thermal spin pumping from detailed spectra remains a formidable challenge Shen (2019). Additionally, conduction electrons interact differently with localized spin and orbital angular momentum Kondo (1962). To focus on competition of antiferromagnetic magnon due to ferrimagnetically coupled spins, we restrict our discussion to gadolinium iron garnet (GdIG), which exhibits zero orbital angular momentum.
Statistical mechanics of GdIG. A model that provides considerable insight into thermal spin pumping in GdIG is the two-magnetic-lattice model Cahaya (2022); Cahaya et al. (2022). In this model, the magnon spectrum of GdIG is determined by two magnetic lattices: $S^{\mathrm{Gd}}$, representing the total angular momentum of three Gd3+ ions, and $S^{\mathrm{Fe}}$, representing the spin angular momentum of Fe3+ ions. This model has been shown to describe both its low-lying spin wave spectrum Tinkham (1961) and room-temperature thermal spin pumping Cahaya (2022).
@@
$\mathcal{H}=$ $-\frac{I}{2}\sum_{n}\textbf{S}^{\mathrm{Fe}}_{j}\cdot\textbf{S}^{\mathrm{Fe}}_{j+1}+\lambda\sum_{j}\textbf{S}^{\mathrm{Gd}}_{j}\cdot\textbf{S}^{\mathrm{Fe}}_{n},$ (9)
where the subscript $n$ indexes the Fe3+ chain. The Fe-Gd coupling induces a paramagnetic response of $S^{\mathrm{Gd}}$ to the molecular field generated by $S^{\mathrm{Fe}}$. Gd-Gd coupling is neglected, as validated by the dominance of Fe-Fe coupling in the Curie temperature $T_{C}=IS^{\mathrm{Fe}}(S^{\mathrm{Fe}}+1)/(3k_{B})+\mathcal{O}\left(\lambda^{2}/I\right)$, Gilleo (1980) obtained from the statistical average:
@@
where $B_{J}(x)$ is the Brillouin function. Fig. 2c illustrates the temperature dependence of the spins. A compensation temperature occurs at Cahaya (2022)
@@
Figure 2: (a) Exchange coupling between spins of Fe3+, occupying tetrahedral and octahedral sites, and Gd3+, occupying dodecahedral sites, in Gd3Fe5555O1212{12}12 (GdIG) can be approximated by (b) the two-spin model with $\mathrm{S}^{\mathrm{Fe}}$ and $\mathrm{S}^{\mathrm{Gd}}$. (c) Temperature dependence of the $z$-component of spins in GdIG. The total spin is zero at the compensation temperature $T_{B}$.
Spin-wave spectrum. The interaction leads to the following Landau–Lifshitz equations of motion for the spins
$\frac{d\textbf{S}^{\mathrm{Fe}}_{j}}{dt}=$ $\textbf{S}^{\mathrm{Fe}}_{n}\times\left(\frac{I}{2}\left(\textbf{S}^{\mathrm{Fe}}_{j+1}+\textbf{S}^{\mathrm{Fe}}_{n-1}\right)-\lambda\textbf{S}^{\mathrm{Gd}}_{j}\right)$
$\frac{d\textbf{S}^{\mathrm{Gd}}_{j}}{dt}=$ $-\lambda\textbf{S}^{\mathrm{Gd}}_{j}\times\textbf{S}^{\mathrm{Fe}}_{j},$ (12)
@@
$i\frac{d}{dt}\left[\begin{array}[]{c}S^{\mathrm{Fe}}_{+}\\ S^{\mathrm{Gd}}_{+}\end{array}\right]=W\left[\begin{array}[]{c}S^{\mathrm{Fe}}_{+}\\ S^{\mathrm{Gd}}_{+}\end{array}\right].$ (17)
$W=\left[\begin{array}[]{cc}IS^{\mathrm{Fe}}_{z}(1-\cos ka)-\lambda S^{\mathrm{Gd}}_{z}&\lambda S^{\mathrm{Fe}}_{z}\\ \lambda S^{\mathrm{Gd}}_{z}&-\lambda S^{\mathrm{Fe}}_{z}\end{array}\right].$ (20)
For small $k$, the eigen frequencies are
$\omega_{1}=$ $\frac{S^{\mathrm{Fe}}_{z}D_{\mathrm{YIG}}k^{2}}{S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}},$
$\omega_{2}=$ $-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+\frac{S^{\mathrm{Gd}}_{z}D_{\mathrm{YIG}}k^{2}}{S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}},$ (21)
where $D_{0}=Ia^{2}S^{\mathrm{Fe}}_{z}/2$ is the spin wave stiffness of YIG. The eigenfrequencies $\omega_{R}=\mathrm{max}\left(\omega_{1},\omega_{2}\right)$ and $\omega_{L}=\mathrm{min}\left(\omega_{1},\omega_{2}\right)$ are associated with right-handed (R) and left-handed (L) magnons, respectively. Their precession directions are illustrated in Fig. 1. Since $S^{\mathrm{Gd}}_{z}$ varies according to Eq. 10, the spectrum changes with temperatures, as illustrated in Fig. 3.
Figure 3: Spin-wave spectrum of GdIG at temperatures (a) lower than, (b) equal to, and (c) higher than the compensation temperature $T_{B}$. Positive frequency ($\omega_{R}$) is associated with right-handed magnon. On the other hand, negative frequency ($\omega_{L}$), with a minimum frequency $\omega_{\mathrm{min}}\simeq-\lambda S^{\mathrm{Fe}}_{z}$, is associated with right-handed magnonboth magnon branches.

### v2

# Spin current compensation from competing magnon modes in ferrimagnets

@@ Abstract
We investigate thermal spin pumping in gadolinium iron garnet (GdIG), focusing on the mode-resolved dynamics of antiferromagnetic magnons and their impact on spin and heat transport. Antiferromagnets support both right-handed and left-handed magnon modes, which we treat as positive and negative frequency branches, analogous to electrons and holes in semiconductors. Using a two-sublattice model with a minimal exchange interaction scheme, we derive the temperature-dependent spin-wave spectrum and evaluate the associated thermal spin pumping coefficients. Our analysis reveals that the competition between left- and right-handed modes gives rise to a compensation temperature, where the net thermally generated spin current vanishes. Importantly, we show that this compensation point does not necessarily coincide with the crossing of magnon dispersion branches. While previous research considers a detailed microscopic model including all magnetic sublattices and exchange couplings, our approach demonstrates that key features of mode-resolved spin transport can be captured by a simplified and analytically transparent model. These findings advance the understanding of spin-caloritronic phenomena in ferrimagnets and offer new perspectives for the design of chiral magnon-based spintronic devices.
[…]
1 Introduction
The exploration of fundamental quasiparticles has driven significant advancements in our understanding of quantum mechanics and condensed matter physics. Among these quasiparticles, magnons, the quanta of spin waves, play a pivotal role in both theoretical frameworks and practical applications in spintronics and quantum information science [1, 2]. In particular, antiferromagnetic magnons, which arise in systems with alternating spin orientations, exhibit rich physical properties due to their multiple propagation modes and interactions.
Building on this concept, researchers have introduced magnons with distinct handedness, referring to the direction of spin precession: left-handed and right-handed. This property emerges due to the underlying magnetic symmetry and adds a new degree of freedom in magnonic systems, enabling chirality-based manipulation of spin waves [3]. Thermal spin pumping has become a promising technique for detecting and utilizing these chiral magnon modes [4]. By converting thermally excited spin waves into measurable electrical signals, this method provides an indirect but powerful probe of the magnonic spectrum.
Rare-earth iron garnets (RIGs) are particularly well-suited for studying such effects due to their complex magnetic structures and ultra-low magnetic damping [5]. The three magnetic sublattices — tetrahedral Fe3+, octahedral Fe3+, and dodecahedral R3+ sites — give rise to multiple magnonic branches [6, 7]. Tuning the rare-earth ion content enables precise control of the spin-wave modes [8]. Moreover, strong spin-orbit coupling effects, such as those from the Dzyaloshinskii-Moriya interaction and the inverse Faraday effect, further enhance their potential for spintronic applications [9]. A particularly intriguing feature in thermal spin pumping in ferrimagnets is the presence of two compensation temperatures. The first corresponds to the compensation of net magnetic moment, which is well-established. The second, at a lower temperature, has been proposed to result from the competition between left-handed and right-handed magnons. In Ref.[10], this lower compensation point was observed to coincide with the crossing of the $|\boldsymbol{\omega}_{R}|$ and $|\boldsymbol{\omega}_{L}|$ spectra, suggesting a connection between spectral symmetry and thermal spin current reversal. However, the separate contributions of left- and right-handed modes to thermal spin pumping have not yet been quantitatively evaluated.
In this manuscript, we address this gap by explicitly calculating the contributions of each mode using a two-magnetic-lattice model, allowing mode-resolved analysis of thermal spin currents. Our model reveals that the lower compensation temperature does not align with the spectral crossing of $|\boldsymbol{\omega}|$, and conversely, that a spectral crossing can occur even when thermal spin pumping remains finite. These findings question the assumption that compensation behavior is directly linked to spectral symmetry and highlight the importance of a detailed, mode-specific approach.
@@ 1 Introduction
To interpret these results, we treat right-handed and left-handed magnon modes as positive and negative frequency branches, respectively. This treatment draws an analogy to electronic transport in semiconductors, where electrons and holes contribute oppositely to current due to their energy band positions. By applying this analogy to spin transport, we develop a formalism in which the net thermal spin current arises from the imbalance between left- and right-handed magnon populations, governed by their respective dispersion relations and thermal distributions.
While Ref. [10] employs a comprehensive microscopic model of GdIG that includes detailed exchange interactions among all magnetic sublattices, our approach is deliberately simplified. We adopt a two-magnetic-lattice model incorporating only the essential Fe–Fe and Fe–Gd exchange interactions. This reduced model is analytically tractable and transparent, yet sufficiently rich to capture key features of the mode-resolved magnon dynamics, including the emergence of compensation temperatures and the role of spectral symmetry in spin transport.
The remainder of this paper is organized as follows: In Section 2, we outline the formalism for calculating spin and heat currents based on the spin-mixing conductance and magnon dynamics, and derive analytic expressions for the spin transport coefficients $\mathcal{L}_{0}$ and $\mathcal{L}_{1}$. In Section 3, we apply a two-sublattice model of GdIG to compute the temperature-dependent magnon spectrum and investigate the contributions of right- and left-handed magnons to thermal spin pumping. We identify the compensation conditions and provide a quantitative analysis of mode-resolved spin transport. Finally, we summarize our conclusions and discuss future perspectives in Section 4.
2 Methods
In thermal spin pumping, spin $J_{s}$ and heat $J_{Q}$ currents are transferred from a magnetic material, such as rare-earth iron garnet, to a non-magnetic metal, such as Pt. This spin current is then converted into an electrical signal via the inverse spin Hall effect. At the interface, $J_{s}$ and $J_{Q}$ coupled and can be expressed through a linear matrix relation involving the driving forces: spin accumulation $\boldsymbol{\mu}_{s}$ and temperature difference $\Delta T$, analogous to the thermoelectric effect:
$\left[\begin{array}[]{c}J_{s}\\ J_{Q}\end{array}\right]=\left[\begin{array}[]{cc}\mathcal{L}_{0}&\mathcal{L}_{1}\\ \mathcal{L}_{1}&\mathcal{L}_{2}\end{array}\right]\left[\begin{array}[]{c}-\hat{\textbf{m}}\cdot\boldsymbol{\mu}_{s}\\ -\Delta T/T\end{array}\right].$ (1)
Here, $\hat{\textbf{m}}=\hat{z}S_{z}/|S_{z}|$ is the average direction of the magnetization. $\mathcal{L}_{i}$ is analogous to thermoelectric transport coefficients. $\mathcal{L}_{0}$ represents the interface spin injection conductance, while $\mathcal{L}_{2}$ denotes the heat conductance. Since the heat conductance at the interface is generally dominated by Kapitza resistance [11], our discussion is focused on $\mathcal{L}_{0}$ and $\mathcal{L}_{1}$. The ratios $\mathcal{L}_{1}/\mathcal{L}_{0}$ and $\mathcal{L}_{1}/(\mathcal{L}_{0}T)$ correspond to the spin Peltier and spin Seebeck coefficients, respectively [11].
@@ 2 Methods
$J_{\mathrm{fl}}=-\gamma\left<\textbf{S}\times\textbf{h}^{\prime}\right>\cdot\hat{z}$ is spin backflow due to thermal noise. $J_{\mathrm{fl}}=-J_{\mathrm{sp}}(T-\Delta T)/T$ has been shown to balances $J_{\mathrm{sp}}$ when there is no temperature difference across interface [12].
@@
where $G$ is the spin-mixing conductance[13].
@@
$\mathcal{L}_{0}$ $=$ $-\frac{\partial J_{s}}{\partial\hat{\textbf{m}}\cdot\boldsymbol{\mu}_{s}}=\frac{G|S_{z}|}{2S_{z}}\left<S_{-}S_{+}+S_{+}S_{-}\right>$ (6)
$\simeq$ $2G|S_{z}|\left<a^{\dagger}a\right>.$
@@
$\mathcal{L}_{0}$ $=$ $2GS_{z}\int_{\omega_{0}}^{\infty}\frac{\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}$ (7)
$=$ $\frac{GS_{z}}{2}\left(\frac{k_{B}T}{\pi D}\right)^{\frac{3}{2}}\mathrm{Li}_{\frac{3}{2}}\left(e^{-\beta\omega_{0}}\right),$
@@
$=$ $2GS_{z}\int_{0}^{\infty}\frac{\omega\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}$
$=$ $\frac{3GS_{z}\left(k_{B}T\right)^{\frac{5}{2}}}{4\left(\pi D\right)^{\frac{3}{2}}}\mathrm{Li}_{\frac{5}{2}}\left(e^{-\beta\omega_{0}}\right).$
In YIG, the magnon gap $\omega_{0}=\gamma H$ can be non zero due to Zeeman effect [14]. Since $\lim_{x\to 0}\mathrm{Li}_{n}(x)\simeq x$, the magnetic field $H$ exponentially suppress thermal spin pumping [15] $\mathcal{L}_{1}\propto e^{-\gamma H/k_{B}T}.$ The magnetization of YIG originates from five Fe3+ ions distributed across two magnetic lattices. Two of these ions occupy tetrahedral Fe3+ sites, while three are in octahedral sites. Due to their antiferromagnetic coupling, only one Fe3+ spin contributes [16], as illustrated in Fig. 2a. When yttrium in YIG is substituted by lanthanides, R3+ occupies the dodecahedral site, forming an additional magnetic lattice. This substitution leads to a more complex spin-wave spectrum compared to that of YIG. While significant computational efforts are underway to unravel these spectra [17, 10], determining thermal spin pumping from detailed spectra remains a formidable challenge [18]. Additionally, conduction electrons interact differently with localized spin and orbital angular momentum [19]. To focus on competition of antiferromagnetic magnon due to ferrimagnetically coupled spins, we restrict our discussion to gadolinium iron garnet (GdIG), which exhibits zero orbital angular momentum.
2.1 Statistical mechanics of GdIG.
A model that provides considerable insight into thermal spin pumping in GdIG is the two-magnetic-lattice model [20, 21]. In this model, the magnon spectrum of GdIG is determined by two magnetic lattices: $S^{\mathrm{Gd}}$, representing the total angular momentum of three Gd3+ ions, and $S^{\mathrm{Fe}}$, representing the spin angular momentum of Fe3+ ions. This model has been shown to describe both its low-lying spin wave spectrum [22] and room-temperature thermal spin pumping [20].
Figure 2: (a) Exchange coupling between spins of Fe3+, occupying tetrahedral and octahedral sites, and Gd3+, occupying dodecahedral sites, in Gd3Fe5O1212{12}12 (GdIG) can be approximated by (b) the two-spin model with $\mathrm{S}^{\mathrm{Fe}}$ and $\mathrm{S}^{\mathrm{Gd}}$. (c) Temperature dependence of the $z$-component of spins in GdIG is in agreement to Ref. [23]. The total spin is zero at the compensation temperature $T_{B}$.
@@
$\mathcal{H}=$ $-\frac{I}{2}\sum_{j}\textbf{S}^{\mathrm{Fe}}_{j}\cdot\textbf{S}^{\mathrm{Fe}}_{j+1}+\lambda\sum_{j}\textbf{S}^{\mathrm{Gd}}_{j}\cdot\textbf{S}^{\mathrm{Fe}}_{j},$ (9)
where the subscript $n$ indexes the Fe3+ chain. The Fe-Gd coupling induces a paramagnetic response of $S^{\mathrm{Gd}}$ to the molecular field generated by $S^{\mathrm{Fe}}$. Gd-Gd coupling is neglected, as validated by the dominance of Fe-Fe coupling in the Curie temperature $T_{C}=IS^{\mathrm{Fe}}(S^{\mathrm{Fe}}+1)/(3k_{B})+\mathcal{O}\left(\lambda^{2}/I\right)$, as derived in the A. The effect of Gd-Gd exchange, which is order of magnitudes smaller than $\lambda$ [24], is also estimated in the Appendix. Fig. 2c illustrates the temperature dependence of the spins, obtained from the statistical average:
@@
where $B_{J}(x)$ is the Brillouin function. A compensation temperature occurs at [20]
@@
For GdIG, $T_{B}\simeq 0.5T_{C}$, $\lambda\simeq 0.1I$.
Compared to the detailed microscopic approach taken in Ref. [10], which models all magnetic sublattices and their exchange interactions explicitly, our two-magnetic-lattice model is highly simplified. It includes only the dominant Fe–Fe and Fe–Gd isotropic exchange couplings and neglects orbital contributions, dipolar effects, and anisotropies. This simplification allows for an analytically tractable framework that nonetheless reproduces key features of the magnonic structure and spin transport in GdIG, including the temperature dependence of compensation behavior and mode competition.
[…]
The interaction leads to the following Landau–Lifshitz equations of motion for the spins
$\frac{d\textbf{S}^{\mathrm{Fe}}_{j}}{dt}$ $=$ $\textbf{S}^{\mathrm{Fe}}_{j}\times\left(\frac{I}{2}\left(\textbf{S}^{\mathrm{Fe}}_{j+1}+\textbf{S}^{\mathrm{Fe}}_{j-1}\right)-\lambda\textbf{S}^{\mathrm{Gd}}_{j}\right)$
$\frac{d\textbf{S}^{\mathrm{Gd}}_{j}}{dt}$ $=$ $-\lambda\textbf{S}^{\mathrm{Gd}}_{j}\times\textbf{S}^{\mathrm{Fe}}_{j}.$ (12)
@@
$i\frac{d}{dt}\left[\begin{array}[]{c}S^{\mathrm{Fe}}_{+}\\ S^{\mathrm{Gd}}_{+}\end{array}\right]=\left[\begin{array}[]{cc}IS^{\mathrm{Fe}}_{z}(1-\cos ka)-\lambda S^{\mathrm{Gd}}_{z}&\lambda S^{\mathrm{Fe}}_{z}\\ \lambda S^{\mathrm{Gd}}_{z}&-\lambda S^{\mathrm{Fe}}_{z}\end{array}\right]\left[\begin{array}[]{c}S^{\mathrm{Fe}}_{+}\\ S^{\mathrm{Gd}}_{+}\end{array}\right],$
the eigen frequencies are
$\omega=\frac{2D_{\mathrm{YIG}}}{a^{2}}\sin^{2}\frac{ka}{2}-\lambda\frac{S_{z}^{\mathrm{Fe}}+S_{z}^{\mathrm{Gd}}}{2}$
±$\pm\sqrt{\left(\lambda\frac{S_{z}^{\mathrm{Fe}}+S_{z}^{\mathrm{Gd}}}{2}\right)^{2}-\frac{4\lambda D_{\mathrm{YIG}}S_{z}^{\mathrm{Gd}}}{a^{2}}\sin^{2}\frac{ka}{2}+\frac{4D^{2}_{\mathrm{YIG}}}{a^{4}}\sin^{4}\frac{ka}{2}}$
where $D_{\mathrm{YIG}}=Ia^{2}S^{\mathrm{Fe}}_{z}/2$ is the spin wave stiffness of YIG. Since $S^{\mathrm{Gd}}_{z}$ varies according to Eq. 10, the spectrum changes with temperatures, as illustrated in Fig. 3. The positive and negative eigen frequencies are associated with right-handed (R) and left-handed (L) magnons, respectively. Their precession directions are illustrated in Fig. 1.
Figure 3: Spin-wave spectrum of GdIG at temperatures (a,d) lower than, (b,e) equal to, and (c,f) higher than the compensation temperature $T_{B}$. Positive frequency ($\omega_{R}$) is associated with right-handed magnon. On the other hand, negative frequency ($\omega_{L}$), with a minimum frequency $\omega_{\mathrm{min}}\simeq-\lambda S^{\mathrm{Fe}}_{z}$, is associated with right-handed magnon branches.
3 Results and Discussion